# May I Ask a Follow-up Question? Understanding the Benefits of Conversations in Neural Network Explainability

### v1

@@ 4. Results & Discussion
This section presents the experimental results of the objective and subjective evaluations before and after conversational explanations with XAI experts. We also analyze collected conversations and discuss the implications of our study for building dialogue systems to explain static explanations in the future. We report the statistical significance of the results using Student’s T-tests.
4.1. Conversational explanations improve comprehension of static explanations
@@
4.2. Objective performance varies between LIME and Grad-CAM
Grad-CAM has a higher accuracy of objective evaluation compared to LIME. A potential reason might be the inherently intuitive nature of the explanations produced by Grad-CAM compared to LIME. This aligns with prior research on different explanations (Kim et al., 2023), where heatmaps were highlighted as more intuitive. Even though Grad-CAM is more intuitive to understand and participants can intuitively choose the best classification model based on the heatmap in the objective evaluation, participants’ self-reporting comprehension score is just slightly larger than 4 (average understanding). This shows that participants still feel confused about how Grad-CAM works and how it explains models’ predictions, even though they can perform well in the objective evaluation.
This also demonstrates that subjective and objective evaluations measure participants’ understanding of static explanations from complementary aspects. Self-reporting scores can be influenced by personal biases, while the objective evaluation might not capture users’ feelings about understanding. Combining both methods can provide a comprehensive view of participants’ understanding of static explanations.
Discipline Evaluation Timing Objective Understanding (Decision-Making Accuracy) Subjective Understanding Perceived Usefulness Perceived Ease of Use Behavioral Intention Trust
Business before 0.63 $\pm$ 0.36 4.00 $\pm$ 0.88 5.26 $\pm$ 1.02 4.43 $\pm$ 1.07 5.29 $\pm$ 0.75 4.32 $\pm$ 0.96
after 0.77 $\pm$ 0.22 5.00 $\pm$ 0.96 5.91 $\pm$ 0.72 5.29 $\pm$ 0.74 5.75 $\pm$ 0.75 5.00 $\pm$ 0.93
Engineering before 0.53 $\pm$ 0.34 4.50 $\pm$ 1.60 5.21 $\pm$ 1.49 4.67 $\pm$ 0.67 5.19 $\pm$ 1.07 4.16 $\pm$ 1.15
after 0.65 $\pm$ 0.26 5.38 $\pm$ 1.19 6.04 $\pm$ 0.57 5.50 $\pm$ 0.76 5.94 $\pm$ 1.15 5.09 $\pm$ 0.79
Humanity pre-conv 0.67 $\pm$ 0.26 4.13 $\pm$ 1.14 5.46 $\pm$ 1.05 4.55 $\pm$ 1.01 5.57 $\pm$ 0.79 4.35 $\pm$ 0.97
post-conv 0.80 $\pm$ 0.22 5.52 $\pm$ 0.85 6.16 $\pm$ 0.63 5.46 $\pm$ 0.98 6.11 $\pm$ 0.74 5.11 $\pm$ 0.78
Science pre-conv 0.51 $\pm$ 0.28 3.93 $\pm$ 1.16 5.09 $\pm$ 0.80 4.78 $\pm$ 0.9 5.30 $\pm$ 0.82 4.18 $\pm$ 0.84
post-conv 0.60 $\pm$ 0.25 5.47 $\pm$ 0.83 5.91 $\pm$ 0.62 5.47 $\pm$ 0.77 5.93 $\pm$ 0.78 5.02 $\pm$ 0.66
Table 4. Results of objective evaluation (decision-making accuracy), self-reporting comprehension score, acceptance, and trust of explanation for participants from different disciplines groups before and after conversations. Scores are presented as mean $\pm$ standard deviation. We do not observe statistically significant differences among participants from different disciplines.
4.3. Conversational explanations improve acceptance and trust of static explanations
@@
(a) Results before conversational explanations
(b) Results after conversational explanations.
Figure 5. Results of objective evaluation (decision-making accuracy), self-reporting comprehension score, acceptance, and trust of explanation for participants from different disciplines groups before and after conversational explanations. Decision-making accuracy is ranged from 0 to 1 and the rest self-reported scores are from 1 to 7. Figure (a) shows the results before conversational explanations and (b) shows the results after conversational explanations.
4.4. No statistically significant differences observed among participants from different disciplines
To investigate whether participants from different disciplines perform differently, we categorize their disciplines into the following four groups:
  • Business: Business and Accountancy.
  • Engineering: Civil and Environmental engineering, Computer Science, Electrical and Electronics Engineering, Maritime Studies, and Food Science.
  • Humanities: Psychology, Economics, Communication Studies, Linguistics and Multilingual Studies, and Sociology.
  • Science: Biology, Chemistry, Chemical Engineering and Biotechnology, Sport Science & Management, Mathematics, and Medicine.
Table 4 and Figure 5 display the results of participants from these discipline categories. We do not observe statistically significant differences among different discipline groups. We employ the one-way ANOVA to validate our observations. For pre-conversation results, the significance values are as follows: decision-making accuracy ($p=0.36$), self-reporting comprehension score ($p=0.71$), perceived usefulness score ($p=0.75$), perceived ease of use ($p=0.79$), behavioral intention ($p=0.60$), and trust ($p=0.94$). The post-conversation evaluations show significance values of 0.06, 0.39, 0.58, 0.91, 0.63, and 0.97, respectively.
Participants from different disciplines do not seem to benefit from discipline-specific training when trying to understand the explanations. Although one may be tempted to conjecture that people from different disciplines would possess different skills, which may lead to easier or more difficult experiences working with AI, our results do not reveal any such differences. This may suggest that understanding AI and XAI requires highly specialized background knowledge that are not shared by the four broad categories we used in the study. Hence, the duty falls upon the XAI designers to effectively communicate with the users wihout assuming any knowledge background. Conversational explanations may function as one such form of communication.
Question Category
Question Examples
Num
Basic concepts in machine learning
  • $\bullet$
    What is a deep learning model?
  • $\bullet$
    What is the image classification task?
  • $\bullet$
    How does the model know what features to extract?
85
Application, performance, and limitations of machine learning models
  • $\bullet$
    How about the precision of the classification model?
  • $\bullet$
    Where has this swin Transformer classification method been used in practical applications?
  • $\bullet$
    Will the different species of an animal affect the classification model categorizing the animal?
68
Diagram reading
  • $\bullet$
    Are regions colored in red areas that have been identified as containing key features for the animal?
  • $\bullet$
    What are the yellow line spots for (in LIME explanations)?
  • $\bullet$
    What do the red color and blue color mean (in Grad-CAM explanations)?
16
Basic concepts of explanation methods
  • $\bullet$
    What is the explanation model used for?
  • $\bullet$
    Can LIME be used without the internet?
  • $\bullet$
    What are some limitations of the Grad-CAM (LIME) method?
95
Mechanism of explanation methods
  • $\bullet$
    Why does the (LIME) explanation not highlight all the parts of the leopard?
  • $\bullet$
    How LIME model recognize the most important parts for the model prediction?
  • $\bullet$
    Seems like the Classification Model and the Explanation Model are trained separately - how can we be sure that the underlying logic of making a prediction is the same for both models?
91
New explanations
  • $\bullet$
    Can you list other visualization methods?
  • $\bullet$
    Is there anything special about the Grad-CAM (or LIME) method that is different from others?
  • $\bullet$
    What if there are both fishes and humans in an image? How should this image be classified, and can you provide such explanations?
42
Table 5. Overview of Collected Questions. Including categories of questions, examples, and the count of questions in each category. Questions are collected based on conversations between participants and XAI experts about an explanation generated by Grad-CAM or LIME.
@@ 4.5. Analysis of Collected Conversations
We collect 60 free-form conversations between XAI experts and participants from 19 different disciplines. On average, each conversation had 27.4 turns, with each turn comprising approximately 14.4 tokens. By analyzing the users’ questions, we divide them into four categories:
@@
  • Application and performance of machine learning models: Questions about the ability, accuracy, and limitations of machine learning models.
  • Diagram reading: Questions about the explanation diagram generated by Grad-CAM or LIME, e.g., what different color represents in the heatmap.
@@
  • New explanations: Questions that require the generation of new explanations, including from other explanation methods, comparison between the provided explanation and other, currently unavailable explanation methods, or explanations for different predictions.
Based on this categorization, we build a repository for questions that could occur in the conversations. In total, we collect 397 questions from the four different categories. Table 5 contains examples and the number of questions in each category. As observed in Table  5, the questions of participants mainly revolve around basic concepts in machine learning, the fundamentals of explanation methods, and their underlying mechanisms. This trend might be attributed to the multi-disciplinary of the participants. It suggests that many participants may not be familiar with machine learning models and explanation methods, which is aligned with the real application of explanation methods. Therefore, it’s crucial to tailor responses to these questions to help users better understand explanations. Furthermore, we note a marked interest in new explanations. This could indicate that as users become more familiar with provided explanation examples, they exhibit curiosity about alternative explanation methods and how models might behave under specific scenarios. Concurrently, the diagram reading category contains only 16 questions, implying that explanations generated by Grad-CAM and LIME were relatively straightforward and easy to understand. The diverse range of questions sourced from our conversations underscores that static, one-off explanations are often insufficient for users to understand them. Engaging in dialogue can provide more dynamic and tailored explanations to users, hence deepening their understanding of static explanations.
4.6. User questions are unanticipated and diverse
@@
Several participants misunderstand the idea of the heatmap produced by Grad-CAM as depicting literal heat dissipating from objects. They infer that the model uses the temperature of objects to perform classification. In reality, a heatmap is just a metaphor that visualizes numerical values distributed spatially, which refer to the feature importance in our case. This misconception leads to questions about how the heat of objects is measured and why non-living objects are warmer than their environment. Some example utterances from participants include: “So the Grad-cam method basically just refers to the usage of generating a heatmap to capture living matters correct? … based on the parts of the image that generate more heat?” –P36, “basically using heat to predict what is the input right?…how will we know what is the animal or input simply based on heat?” – P47, “if these are pictures, how do they figure out the heat since the animal isn’t generating heat” – P49, “So a heat sensor is not required? A heatmap is automatically generated from each photo and analyzed using the model.” – P52.
@@
A second common misconception is the conflation between the explanation technique and the classification models. Some example user questions include: “is the explanation method what the model uses to classify & predict what the image is supposed to be?” – P6, “swin transformer uses LIME model? … what’s the differences between lime model and swin transformer?” – P8. Furthermore, participants face challenges in understanding certain terms commonly used in AI and XAI, even though these terms are frequently used and understood within academic communities. Many participants asked questions about basic concepts in machine learning, such as: “what is explanation method?” – P7, “how do you classify the image?” – P17, “what is the algorithm? does it means lime? what are deep neural networks?” – P32, “How would you explain the term ”perturbations of images” to a five-year-old?” – P46.
@@
The observations from the interactions between XAI experts and layperson users demonstrate the importance of conversations for users to understand static explanations as they bridge the knowledge gap between the two groups. Conversations can reveal the specific areas of misunderstanding, assumptions users might make, and knowledge they might lack. Hence, conversational explanations may help the AI system communicate with and bring genuine understanding to the users.
@@ 4.7. Implications for building dialogue systems to explain static explanations
The emergence of advanced conversational agents (Zhang et al., 2022; Ni et al., 2023; Shuster et al., 2022), especially knowledge-based question-answering (Lan et al., 2021; Luo et al., 2023; Zhang et al., 2023) powered by large language models (Touvron et al., 2023; Ouyang et al., 2022; Zhao et al., 2023) paves the way to build dialogue systems that can explain model decisions and explanations in conversations. Below, we explore the implications of our study for designing such dialogue systems.
@@
  • Extensive Coverage of AI and XAI Concepts. As observed in our study, there are a large portion of questions are the core concepts of machine learning models and explanation methods. To answer those questions, dialogue systems need to be trained on a comprehensive corpus encompassing AI and XAI concepts. Besides, in our study, participants also are curious about the applications, performances, and limitations of machine learning models and explanation methods. Therefore, besides answering theoretical questions, dialogue systems also should be able to link them to real-world applications and limitations.
@@
  • Capability to Generate New Explanations. As an improved understanding of the provided explanations, participants in our study exhibit curiosity about alternative explanation methods and different explanation examples. Dialogue systems should be designed to provide new explanations to users. They should not only retrieve stored explanations but also construct new ones based on the evolving understanding of the user’s question. For instance, if a user is curious about how changing a feature would affect a model’s output, the system should dynamically generate a new explanation, elucidating the potential impact.
@@
  • Ability to Interpret Diagrams. A significant portion of AI and XAI explanations often comes in the form of diagrams (Selvaraju et al., 2017; Ribeiro et al., 2016), such as heatmaps from Grad-CAM or feature importance via LIME. While our study observed fewer questions about diagrams, those questions are essential for users to understand explanations. Therefore, future dialogue systems should have visual processing capabilities, understanding and interpreting diagrams contextually. For instance, they should be able to recognize colors, patterns, and other graphical elements in heatmaps or charts and relate them to users’ questions. The recent development in multimodal large language models (Zhu et al., 2023; Driess et al., 2023; Gong et al., 2023) is a promising direction to achieve this goal.
5. Limitation and Future Work
@@
Our study exploits the effects of free-form conversations that accompany and clarify static explanations. Despite the insights gained, there are several limitations that should be acknowledged. First, the static explanations used in our study are limited. Our experiments focused on feature attribution explanation methods; we selected LIME and Grad-CAM, which are representative of gradient-based methods and surrogate methods respectively. The applicability of our findings to other explanation methods, such as example-based explanation methods, remains an open question. Second, as our main objective was to discern the effects of free-form conversational explanations, we did not delve into the comparative performance of different explanation methods. In our experiments, we intentionally selected explanation examples where the best classification model yielded more reasonable explanations. The explanation examples discussed by participants and XAI experts were chosen such that they reasonably explain the predictions of the classification model. Future work would be to extend these conversations to include explanations that might be less reliable. Third, we investigate how conversations can affect user trust in explanations of AI systems. It’s noteworthy that prior studies have highlighted that humans may place undue trust in AI models even when system decisions are flawed (Wang and Yin, 2021; Zhang et al., 2020). Within the scope of the current study, we do not specifically evaluate if this trust exhibited by users is misplaced or justified. Fourth, we explore how conversations foster user trust in explanations in our study. Nevertheless, previous studies (Wang and Yin, 2021; Zhang et al., 2020) have shown that humans may trust AI models even if they make wrong decisions. We do not explore whether users’ trust in our study is misplaced trust or not, which can be extended in the future. Finally, our research is confined to one (anonymized) geographical region. Cultural nuances may influence the behaviors of users of XAI and how they seek to clarify confusion. Future studies could involve recruiting participants from diverse countries or regions to draw a more comprehensive conclusion.

### v2

@@ 4. Results & Discussion
Table 3 tabulates the mean and standard deviation (SD) for all the measures. As explanation methods (LIME vs. Grad-CAM) and group (experimental vs. control) are between-subjects variables and time (before vs. after) is a within-subject variable, we conduct a three-way Analysis of Variance (ANOVAs).
4.1. Effects of explanations on objective decision accuracy and subjective measures
@@
Figure 4. Objective decision accuracy for different groups before and after conditions.
[…]
Figure 5. Subjective understanding score for (a) LIME and (b) Grad-CAM before and after conditions.
@@
Figure 6. Participants’ self-report usefulness score for (a) LIME and (b) Grad-CAM before and after conditions.
@@
Figure 7. Participants’ self-report ease of use score for (a) LIME and (b) Grad-CAM before and after conditions.
[…]
Figure 8. Participants’ self-report behavioral intention score for (a) LIME and (b) Grad-CAM before and after conditions.
[…]
The boost in usefulness, ease of use, and behavioral intention for the experimental group can be attributed to the increased understanding of static explanations. Prior to the expert interactions, participants might have had limited knowledge or even misconceptions about the explanation methods. Experiment results show that participants gain a clearer understanding of how the XAI methods function, after the participants’ questions are addressed in the conversations. Consequently, they report perceiving the static explanations as more useful and easier to use, and report higher inclination to use the static explanations in future tasks.
The perceived usefulness, ease of use, and behavioral intention of the control group all decrease after reading static explanations for a longer time. This trend suggests a decreased willingness to utilize explanations in future scenarios. The reluctance may be attributed to the frustration the control group faced in attempting to comprehend the static explanations on their own. Research by Carolin Ebermann and Weibelzahl (2023) on the impact of cognitive fit and misfit in the acceptance of AI system usage highlights this phenomenon. They found that users experiencing a cognitive misfit with the AI system often report negative moods, which in turn, reduce their perceived usefulness, ease of use, and behavioral intention of the AI systems. The contrary results of the control group and the experimental group also underscore the importance and effectiveness of conversations in enhancing user behavioral intentions of static explanations.
Figure 9. Participants’ trust for (a) LIME and (b) Grad-CAM before and after conditions.
[…]
@@ 4.2. Analysis of Collected Conversations
We collect 60 free-form conversations between XAI experts and participants from 4 different discipline groups. On average, each conversation had 27.4 turns, with each turn comprising approximately 14.4 tokens. By analyzing the users’ questions, we divide them into six categories:
@@
  • Application and performance of machine learning models: Questions about the ability, accuracy, and limitations of machine learning.
  • Diagram reading: Questions about the explanation diagram generated by Grad-CAM or LIME, e.g., what different colors represent in the heatmap.
@@
  • Other explanations: Questions that require the generation of other types of explanations on the current predictions, explanations for different predictions, or comparisons between the provided explanation and other explanation methods.
Based on this categorization, we build a repository for questions that could occur in the conversations. In total, we collected 397 questions from the four different categories. Table 4 contains examples and the number of questions in each category. As observed in Table 4, the questions of participants mainly revolve around basic concepts in machine learning, the fundamentals of explanation methods, and their underlying mechanisms. This trend might be attributed to the multi-disciplinarity of the participants. It suggests that many participants may not be familiar with machine learning models and explanation methods, which is aligned with the real application of explanation methods. Therefore, it’s crucial to tailor responses to these questions to help users better understand explanations. Furthermore, we note a marked interest in new explanations. This could indicate that as users become more familiar with provided explanation examples, they exhibit curiosity about alternative explanation methods and how models might behave under specific scenarios. Concurrently, the diagram reading category contains only 16 questions, implying that explanations generated by Grad-CAM and LIME were relatively straightforward and easy to understand. The diverse range of questions sourced from our conversations underscores that static, one-off explanations are often insufficient for users to understand them. Engaging in dialogue can provide more dynamic and tailored explanations to users, hence deepening their understanding of static explanations.
@@
Several participants misunderstood the idea of the heatmap produced by Grad-CAM as depicting literal heat dissipating from objects. They infer that the model uses the temperature of objects to perform classification. In reality, a heatmap is just a metaphor that visualizes numerical values distributed spatially, which refers to the feature importance in our case. This misconception leads to questions about how the heat of objects is measured and why non-living objects are warmer than their environment. Some example utterances from participants include: “So the Grad-cam method basically just refers to the usage of generating a heatmap to capture living matters correct? … based on the parts of the image that generate more heat?” –P36, “basically using heat to predict what is the input right?…how will we know what is the animal or input simply based on heat?” – P47, “if these are pictures, how do they figure out the heat since the animal isn’t generating heat” – P49, “So a heat sensor is not required? A heatmap is automatically generated from each photo and analyzed using the model.” – P52.
@@
A second common misconception is the conflation between the post-hoc explanation technique and the classification models. Some example user questions include: “is the explanation method what the model uses to classify & predict what the image is supposed to be?” – P6, “Swin transformer uses LIME model? … what are the differences between lime model and Swin transformer?” – P8. Furthermore, participants face challenges in understanding certain terms commonly used in AI and XAI, even though these terms are frequently used and understood within academic communities. Many participants asked questions about basic concepts in machine learning, such as: “what is the explanation method?” – P7, “how do you classify the image?” – P17, “what is the algorithm? does it mean lime? what are deep neural networks?” – P32, “How would you explain the term ”perturbations of images” to a five-year-old?” – P46.
@@
The observations from the interactions between XAI experts and layperson users demonstrate the importance of conversations for users to understand static explanations as they bridge the knowledge gap between the two groups. Conversations can reveal the specific areas of misunderstanding, such as incorrect implicit assumptions the users make and knowledge they lack. Hence, conversational explanations may help the AI system communicate with and bring genuine understanding to the users.
Table 4. Overview of Collected Questions. Including categories of questions, examples, and the count of questions in each category.
Question Category Question Examples Num
Basic concepts in machine learning $\bullet$ What is a deep learning model? $\bullet$ What is the image classification task? $\bullet$ How does the model know what features to extract? 85
Application, performance, and limitations of machine learning models $\bullet$ How about the precision of the classification model? $\bullet$ Where has this Swin Transformer classification method been used in practical applications? $\bullet$ Will the different species of an animal affect the classification model categorizing the animal? 68
Diagram reading $\bullet$ Are regions colored in red areas that have been identified as containing key features for the animal? $\bullet$ What are the yellow line spots for (in LIME explanations)? $\bullet$ What do the red and blue colors mean (in Grad-CAM explanations)? 16
Basic concepts of explanation methods $\bullet$ What is the explanation model used for? $\bullet$ Can LIME be used without the internet? $\bullet$ What are some limitations of the Grad-CAM (LIME) method? 95
Mechanism of explanation methods $\bullet$ Why does the (LIME) explanation not highlight all the parts of the leopard? $\bullet$ How LIME model recognize the most important parts for the model prediction? $\bullet$ Seems like the Classification Model and the Explanation Model are trained separately - how can we be sure that the underlying logic of making a prediction is the same for both models? 91
Other explanations $\bullet$ Can you list other visualization methods? $\bullet$ Is there anything special about the Grad-CAM (or LIME) method that is different from others? $\bullet$ What if there are both fishes and humans in an image? How should this image be classified, and can you provide such explanations? 42
@@ 4.3. Implications for building dialogue systems to explain static explanations
The emergence of advanced conversational agents (Zhang et al., 2022; Ni et al., 2023; Shuster et al., 2022), especially knowledge-based question-answering (Lan et al., 2021; Luo et al., 2023; Zhang et al., 2023) powered by large language models (Touvron et al., 2023; Ouyang et al., 2022; Zhao et al., 2023) paves the way toward conversational agents that can explain model decisions and discuss static explanations. Our study suggests the following desiderata for such agents.
@@
  • Extensive knowledge of AI and XAI. As observed in our study, a large portion of user questions are related to core concepts of machine learning models and explanation methods. To answer those questions, conversational agents need to be trained on a comprehensive corpus encompassing AI and XAI concepts. Besides, in our study, participants also are curious about the applications, performances, and limitations of machine learning models and explanation methods. Therefore, besides answering abstract questions, dialogue systems also should relate them to real-world applications and limitations.
@@
  • Capability to generate new explanations as needed. As an improved understanding of the provided explanations, participants in our study exhibit curiosity about alternative explanation methods and explaining different predictions. Dialogue systems should provide new explanations to users when requested. For instance, if a user is curious about how changing a feature would affect the model output, the system should generate a new explanation with the new feature, which showcases the effect.
@@
  • Capability to interpret scientific diagrams and visualizations. A significant portion of AI and XAI explanations often comes in the form of diagrams (Selvaraju et al., 2017; Ribeiro et al., 2016), such as heatmaps or feature importance visualizations. Our study reveals that users have questions related to understanding these diagrams. Answering these questions usually requires an understanding of specific regions of the diagrams, such as answering what parts of the object are highlighted by the yellow line in LIME explanations. Therefore, future dialogue systems should have visual processing capabilities, understanding and interpreting diagrams contextually. For instance, they should be able to recognize colors, patterns, and other graphical elements in heatmaps or charts and relate them to users’ questions. The recent development in multimodal large language models (Zhu et al., 2024; Driess et al., 2023; Gong et al., 2023) is a promising direction to achieve this goal.
4.4. Limitations
@@
Despite the insights gained, there are several limitations that should be acknowledged. First, the static explanations used in our study are limited. Our experiments focused on feature attribution explanation methods. The applicability of our findings to other explanation methods, such as example-based explanation methods, remains an open question. Second, as our main objective was to discern the effects of free-form conversational explanations, we did not delve into the comparative performance of different explanation methods. In our experiments, we intentionally selected explanation examples where the best classification model yielded the most reasonable explanations. The explanation examples discussed by participants and XAI experts were chosen such that they reasonably explain the predictions of the classification model. Future work would be to extend these conversations to include explanations that might be less reliable. Third, we explore how conversations foster user trust in explanations in our study. Nevertheless, previous studies (Wang and Yin, 2021; Zhang et al., 2020; Ha and Kim, 2023) have shown that humans may trust AI models even if they make wrong decisions. We do not explore whether users’ trust in our study is misplaced, which we leave for future work. Fourth, we use AI to classify the images. Previous studies (Bankins et al., 2022; Formosa et al., 2022) found that participants favor humans over AI decision-makers when their decisions directly affect participant welfare. In our study, AI decisions do not directly affect participant welfare. We also did not investigate if the participants preferred conversations with humans or AI chatbots or if their trust in the explanations was affected by that variable. Finally, our research is confined to one geographical region and includes only students and staff from the university. Factors such as cultural backgrounds and age-related differences could potentially influence user interactions with XAI and how they seek to clarify confusion. Future studies could involve recruiting participants from diverse countries, regions, and age groups.